### v2

@@ 1.1 Shill Bidding in Auctions
Some tricks of the trade, like an auctioneer’s drumming up excitement by acknowledging nonexistent bids only he hears and potential buyers who bid with nearly imperceptible secret signals, have been around for decades. Making up bids, for instance, is known as “bidding off the chandelier” from an era when the grand auction rooms were adorned with ornate lighting. [^4] See Genteel Auction Houses Turning Aggressive, New York Times, April 24, 2000.
@@
We want to maintain a fair marketplace for all our users, and as such, shill bidding is prohibited on eBay…eBay has a number of systems in place to detect and monitor bidding patterns and practices. If we identify any malicious behavior, we’ll take steps to prevent it. [^6] See https://www.ebay.com/help/policies/selling-policies/selling-practices-policy/shill-bidding-policy?id=4353.
@@
The Sellers post a Buy Now price 3–4 times the actual cost of the item. Then they place the item on an auction at \$0.01. This to get as many views as possible. The shill comes in shortly after the auction starts and …is there to prevent the item from being sold below their profit margin. [^7] See https://community.ebay.com/t5/Buying/My-experience-with-Shill-bidders/td-p/30402514.
@@
Shill-Proof Auctions. Much of auction theory to date encourages truthful bidding through careful auction design, while punting on challenges like seller deviations and collusion via appeal to ummodeled concepts such as the out-of-mechanism enforcement of rules. Anecdotes about eBay and other online platforms suggest that such methods are only partially effective at deterring seller deviations. Thus, it makes sense to ask: To what extent can these deviations instead be disincentivized through an auction’s design?
@@
The goal of this paper is to understand which auction formats are “shill-proof” in the sense that a seller cannot profit through the submission of shill bids. For private-value auctions, which is our focus here, the reader might well wonder why shill bids matter at all—assuming that the choice of reserve price doesn’t affect participation (as it does in the eBay example), isn’t a shill bid the same thing as a reserve price?
@@
The answer depends on when the seller has an opportunity to shill and the information available to them at that time. For example, consider an English auction in which the seller also participates as a shill bidder. Suppose the valuations of the (real) bidders are drawn i.i.d. from a regular distribution $F$ and that the opening bid of the auction is set optimally (for revenue), to the monopoly price $\rho^{*}$ of $F$. As the auction proceeds, with the offered price $p$ starting at $F$ and increasing from there (in increments of $\epsilon$, say), the seller can shill bid at any time. Suppose that the only additional information known to the seller at a given round of the auction is that the remaining bidders are willing to pay at least $p$. Then, the seller asks themself: “now that I know how many bidders are willing to pay at least $p$, do I want to shill and reset the reserve price to $p+\epsilon$?” Under our assumption that $F$ is regular, the answer is “no,” and an expected revenue–maximizing seller will never shill. [^8] Auction theory experts will now immediately recognize that the English auction with an optimally chosen reserve price is not generally shill-proof in this sense when the valuation distribution is not regular., [^9] This doesn’t necessarily mean that the auctioneers at Christie’s are acting suboptimally, as bidders’ valuations in art auctions might be strongly positively correlated—see, for example, Footnote 18 of Milgrom and Weber (1982) for more discussion.
@@
Now suppose that the seller has full knowledge of bidders’ realized valuations. In this scenario, the seller will certainly, in some cases, want to shill in an English auction to push the price up to just below the highest of the bidders’ valuations. Lest this informational assumption—that the seller knows the full valuation profile—seem impossibly demanding, consider the Dutch (descending) auction (with an arbitrary reserve price). Here, any shill bid by the seller terminates the auction immediately, leaving the seller holding the item and earning zero revenue.
@@
  • To what extent are strong and weak shill-proofness compatible with other desirable properties such as optimality, efficiency, ex-post incentive compatiblility, and sealed-bid implementation?
@@ 1.2 Overview of Results
Iterative auction formats like Dutch and English auctions play a central role in our theory, and accordingly we study (real and shill) bidding in the extensive-form game that is induced by a choice of auction format, relying on a framework for extensive-form auction analysis developed by Li (2017) and Akbarpour and Li (2020). We consider single-item auctions with $N$ bidders. A subset of these are shill bidders, which we model as bidders with zero private value for the item and with utility equal to the seller’s revenue. [^10] The prior literature has sometimes modeled shill bidding via an unknown number of bidders, with some subset of the bidders who end up participating in the auction being shills. Our framework is essentially equivalent: we can take $N$ to be large and require 00 to be in the support of the valuation distribution; and a bidder with value 00 is equivalent (in terms of outcomes) to a bidder not arriving. We assume that the private valuations of the non-shill bidders are drawn i.i.d. from a known distribution that is regular (see Definition 2.2). We also assume that shill bidders observe all actions. An auction is then weakly shill-proof (Definition 2.3) if there exists an equilibrium of the induced extensive-form game in which the shill bidders never shill (i.e., always bid their true private value of 00). An auction is strongly shill-proof (Definition 2.4) if, moreover, shill bidders’ equilibrium strategies are ex-post strategies. In our first two results, we focus on public auctions (Definition 3.3), meaning auctions in which every bidder’s action is publicly observable. This is arguably the most natural model for the analysis of typical iterative auctions such as Dutch and English auctions. Note that a static auction like a sealed-bid first-price auction is not public in this sense.
Next, we summarize the main results of this paper; [^11] We focus on single-item auctions, but our uniqueness results a fortiori provide an upper bound on what is possible for multiunit auctions, as well. More broadly, while our analysis is confined to the context of a single auction, we implicitly bound what is achievable in a sequential auction setting. That said, there are further potential incentives for shill bidding in an auction sequence, such as the desire to have a high clearing price for the first auction in order to drive up attention to later ones; it is not clear that any mechanism design could prevent this type of shill bidding incentive in full generality. see also Figure 1.
@@
Strongly Shill-Proof Auctions. Our main result (Theorem 3.4) is a uniqueness result for strongly shill-proof auctions: the Dutch auction (with consistent tie-breaking and monopoly reserve price) is strongly shill-proof and optimal (i.e., maximizes the seller’s expected revenue), and it is the only such auction in the public setting. In particular, strongly shill-proof optimal auctions cannot avoid using a large number of rounds, and they cannot be ex-post incentive compatible (for real bidders). The rough intuition for the proof of this result is that: (i) No matter the information structure, strongly shill-proof auctions must be pay-as-bid (Lemma 3.2); (ii) for any auction format other than a Dutch auction, there exists a history in which some bidder $i$ can indicate that her value is higher than 00 without the auction ending immediately; (iii) optimality in tandem with the public setting then implies that this information effectively induces the auction to revise its reserve price upward or increases perceived competitiveness for the item being sold, which reduces bid shading; and thus (iv) there exist valuations for the bidders such that, if bidder $i$ is a shill bidder, shilling will increase the seller’s revenue.
Weakly Shill-Proof and Efficient Auctions. Our remaining results concern the richer design space of weakly shill-proof auctions (which contains, at the least, both Dutch and English auctions). We start by investigating efficient (and weakly shill-proof) public auctions in which, at equilibrium, the item is always awarded to the (real) bidder with the highest valuation. One example of such an auction is a Dutch auction with a reserve price of 00. Interestingly, English auctions are not examples: with a zero reserve, it is not weakly shill-proof (as shill bidders are motivated to push the price up to the monopoly price). But the Dutch auction is not the only efficient and weakly shill-proof auction: beginning with an English auction at the monopoly price and then, should there be no takers, concluding with a Dutch auction (with no reserve) is another example. [^12] In fact, this auction format closely resembles the Honolulu–Sydney fish auction documented by Hafalir et al. (2023). That this “hybrid” auction format concludes with a Dutch auction is no accident: we prove (in Theorem 3.6) that every robustly weakly shill-proof and efficient auction must conclude with a Dutch auction when all bidders’ valuations fall below the monopoly price $\rho^{*}$ of the distribution. [^13] Here “robustly” means that the auction specification should depend only on the monopoly price $\rho^{*}$ of the valuation distribution, and not on its more fine-grained details. In particular, no ex-post incentive compatible auction can be robustly weakly shill-proof and efficient. It follows that the Dutch auction is the unique prior-independent auction (in the sense of Dhangwatnotai et al. (2015), with no dependence whatsoever on the valuation distribution) that is both efficient and weakly shill-proof (Corollary 3.8).
Weakly Shill-Proof and Ex-Post Incentive Compatible Optimal Auctions. The previous two results imply that ex-post incentive compatible auctions cannot be both strongly shill-proof and optimal, nor can they be (robustly) weakly shill-proof and efficient. The English auction (with an optimal reserve price) is, as we’ve noted, weakly shill-proof, optimal, and ex-post incentive compatible. Is it the unique such auction? Does this combination of properties require a potentially large number of rounds? Our next result (Theorem 4.5) shows that, in general, the answer is no: interrupting an English auction after a sufficiently large number of rounds and closing it out with a second-price auction among the remaining bidders is also weakly shill-proof, optimal, and ex-post incentive compatible. In fact, for any $\varepsilon>0$, we can always find a valuation distribution such that the number of auction rounds that is required in the worst case to maintain weak shill-proofness is an arbitrarily small fraction of the number of rounds needed in the English auction. (The next result implies that the number of rounds cannot be reduced to $1$.)
@@
Proof (Theorem 5.2) (Theorem 4.5)
@@
Proof (Theorem 5.2) (Theorem 3.6)
@@
Proof (Theorem 3.6) (Robustly Unique, Theorem 3.6)
@@
Proof (Theorem 3.6) (Theorem 3.6)
@@
Proof (Theorem 3.6) (Unique, Theorem 3.6)
@@ 1.3 Related Work
While the idea and practice of shill bidding by a seller have long been well known, the auction theory literature on the topic is surprisingly thin. Chakraborty and Kosmopoulou (2004) consider common value auctions and focus on technological barriers (as opposed to auction formats) that can mitigate shill bidding. Lamy (2009) studies shill bidding specifically in English auctions in which bidders’ valuations are affiliated in the sense of Milgrom and Weber (1982), and proves that shill bidding effectively cancels out the effects of affiliation in equilibrium due to real bidders conditioning on bids being fake (see also Izmalkov (2004)). Porter and Shoham (2005) consider a model similar to a second-price auction, motivated by “cheating” by online platforms that can announce a manipulated auction outcome subsequent to collecting all of the bidders’ bids. More recently, a number of works (e.g., Basu et al. (2023); Chung and Shi (2023); Lavi et al. (2022); Roughgarden (2021)) have considered shill bidding in the context of blockchain transaction fee mechanism design, with an emphasis on knapsack auctions that are ex-post incentive compatible, shill-proof, and robust to various forms of collusion. Ausubel and Milgrom (2006) and Day and Milgrom (2008) consider shill bids by bidders in a multi-item auction, who are looking to exploit complementarities to lower their payments in VCG-type mechanisms—as opposed to shill bids by a seller looking to increase revenue, as is the case of this paper. [^15] More broadly there is a literature on sybil resistance referred to as “false-name proofness.” See, e.g., Conitzer et al. (2010) for a reference. Contemporary work by Zeng (2024) also studies shill bidding, but takes a different approach to the problem. While we study ex-interim deterrence against shill bidding in extensive form games, he groups auctions into equivalence classes based on the outcome and primarily focuses on ex-ante deterrence against the seller inserting additional bidders into the auction to increase perceived competition. The results in that paper are complementary to ours: While we show that the dynamics of the auction are important for preventing shill bidding, he shows that, when the shill bidders can increase perceived competition outside of taking actions in the auction, only the posted price auction is non-manipulable. See also Footnote 19.
@@
Our theory of shill-proof auctions is similar in spirit to the theory of credible mechanisms developed by Akbarpour and Li (2020), and leverages their framework for extensive-form auction analysis. That said, shill-proofness differs conceptually from credibility as shill-proofness focuses on the auctioneer’s incentives to insert fake bids whereas credibility focuses on the auctioneer’s incentive to truthfully report the actions of a bidder to other bidders. Further, the results in this paper are also qualitatively different. For example, there are a multitude of credible auctions, but only one strongly shill-proof auction and there are a multitude of strategy-proof and weakly shill-proof auctions, but only one strategy-proof and credible auction (see Section 6.1 for more discussion).

### v3

@@ 1.1 Shill Bidding in Auctions
Some tricks of the trade, like an auctioneer’s drumming up excitement by acknowledging nonexistent bids only he hears and potential buyers who bid with nearly imperceptible secret signals, have been around for decades. Making up bids, for instance, is known as “bidding off the chandelier” from an era when the grand auction rooms were adorned with ornate lighting. [^4] See “Genteel Auction Houses Turning Aggressive,” New York Times, April 24, 2000.
@@
We want to maintain a fair marketplace for all our users, and as such, shill bidding is prohibited on eBay. […] eBay has a number of systems in place to detect and monitor bidding patterns and practices. If we identify any malicious behavior, we’ll take steps to prevent it. [^6] See https://www.ebay.com/help/policies/selling-policies/selling-practices-policy/shill-bidding-policy?id=4353.
@@
The Sellers post a Buy Now price 3–4 times the actual cost of the item. Then they place the item on an auction at \$0.01. This is to get as many views as possible. The shill comes in shortly after the auction starts and …is there to prevent the item from being sold below their profit margin. [^7] See https://community.ebay.com/t5/Buying/My-experience-with-Shill-bidders/td-p/30402514.
@@
Shill-Proof Auctions. Much of auction theory to date encourages truthful bidding through careful auction design, while punting on challenges like seller deviations and collusion via appeal to unmodeled concepts such as the out-of-mechanism enforcement of rules. Anecdotes about eBay and other online platforms suggest that such methods are only partially effective at deterring seller deviations. Thus, it makes sense to ask: To what extent can these deviations instead be disincentivized through an auction’s design?
@@
The goal of this paper is to understand which auction formats are “shill-proof” in the sense that a seller cannot profit through the submission of shill bids. We show that shill bidding can matter even in private-value auctions. The reader might wonder why shill bids can have an impact in the private-values case—assuming that the choice of reserve price does not affect participation (as it does in the eBay example), isn’t a shill bid the same thing as a reserve price?
@@
The answer depends on when the seller has an opportunity to shill and the information available to them at that time. For example, consider an English auction in which the seller also participates via shill bidding. Suppose the valuations of the (real) bidders are private and drawn i.i.d. from a regular distribution $F$ and that the opening bid of the auction is set optimally (for revenue), to the monopoly price $\theta^{\rho^{*}}$ of $F$. As the auction proceeds, with the offered price $p$ starting at $\theta^{\rho^{*}}$ and increasing from there (in increments of $\epsilon$, say), the seller can shill bid at any time. Suppose that the only additional information known to the seller at a given round of the auction is that the remaining bidders are willing to pay at least $p$. Then, the seller asks himself: “now that I know how many bidders are willing to pay at least $p$, do I want to shill and reset the reserve price to $p+\epsilon$?” Under our assumption that $F$ is regular, the answer is “no,” and an expected revenue–maximizing seller will never shill. [^8] Auction theory experts will now immediately recognize that the English auction with an optimally chosen reserve price is not generally shill-proof in this sense when the valuation distribution is not regular or the values are not private. See, for example, Footnote 18 of Milgrom and Weber (1982) for more discussion of the latter point.
@@
Now suppose that the seller has full knowledge of bidders’ realized valuations. In this scenario, the seller will certainly, in some cases, want to shill in an English auction to push the price up to just below the highest of the bidders’ valuations. Lest this informational assumption—that the seller knows the full valuation profile—seem impossibly demanding, consider the Dutch (descending) auction (with an arbitrary reserve price). Here, any shill bid by the seller terminates the auction immediately, leaving the seller holding the item and earning zero revenue. Therefore, even if the seller knows the full valuation profile, he would not want to shill bid.
@@
  • To what extent are strong and weak shill-proofness compatible with other desirable properties such as optimality, efficiency, ex-post incentive compatiblility, and sealed-bid implementation? To what extent is shill-proofness dependent on the bidders’ valuation structure?
@@ 1.2 Overview of Results
Iterative auction formats like Dutch and English auctions play a central role in our theory, and accordingly we study (real and shill) bidding in the extensive-form game that is induced by a choice of auction format, relying on a framework for extensive-form auction analysis developed by Li (2017) and Akbarpour and Li (2020). We consider single-item auctions with $N$ bidders. A subset of these are shill bidders, which we model as bidders with zero private value for the item and with utility equal to the seller’s revenue. [^9] The prior literature has sometimes modeled shill bidding via an unknown number of bidders, with some subset of the bidders who end up participating in the auction being shills. Our framework is essentially equivalent: we can take $N$ to be large and require 0 to be in the support of the valuation distribution; and a bidder with value 0 is equivalent (in terms of outcomes) to a bidder not arriving. For most of the paper, we assume that non-shill bidders have private valuations drawn i.i.d. from a known distribution. In Section 5, we generalize to any affiliated type distribution and any interdependent value function satisfying some curvature assumptions (Assumption 5.1). We also assume that shill bidders observe all actions. An auction is then weakly shill-proof (Definition 2.1) if there exists an equilibrium of the induced extensive-form game in which the shill bidders never shill (i.e., always bid their true private value of 0). An auction is strongly shill-proof (Definition 2.2) if, moreover, shill bidders’ equilibrium strategies are ex-post strategies. In our first result, we focus on public auctions (Definition 3.4), meaning auctions in which every bidder’s action is publicly observable. This is arguably the most natural model for the analysis of typical iterative auctions such as Dutch and English auctions. We then turn to arbitrary information structures to prove our other results.
Next, we summarize the main results of this paper; see also Figure 1.
@@
Strongly Shill-Proof Auctions. Our main result (Theorem 3.6) uniquely characterizes strongly shill-proof auctions: the Dutch auction (with consistent tie-breaking and monopoly reserve price) is strongly shill-proof and optimal (i.e., maximizes the seller’s expected revenue), and is the only such auction in the public setting. In particular, strongly shill-proof optimal auctions cannot avoid using a large number of rounds, and they cannot be ex-post incentive compatible (for real bidders). The rough intuition for the proof of this result is that: (i) No matter the information structure, strongly shill-proof auctions must be pay-as-bid (Lemma 3.1); (ii) for any auction format other than a Dutch auction, there exists a history in which some bidder $i$ can indicate that her value is higher than 0 without the auction ending immediately; (iii) incentive compatibility in tandem with the public setting then implies that this information effectively induces the auction to revise its reserve price upward or increases perceived competitiveness for the item being sold, which reduces bid shading; and thus (iv) there exist valuations for the bidders such that, if bidder $i$ is a shill bidder, shilling will increase the seller’s revenue. [^10] We focus on single-item auctions, but our uniqueness results a fortiori provide an upper bound on what is possible for multiunit auctions, as well. We leave formal study of multi-unit auctions and other settings, such as sequential auctions, to future work.
@@
Weakly Shill-Proof and Efficient Auctions. We then investigate efficient (and weakly shill-proof) public auctions. We prove that a Dutch auction with a reserve price of 0 is the unique prior-independent auction (in the sense of Dhangwatnotai et al. (2015), with no dependence whatsoever on the valuation distribution) that is both efficient and weakly shill-proof (Corollary 4.8). In the regular, IPV case, we show (in Proposition 4.6) that fixing the monopoly price, a prior-independent, weakly shill-proof and efficient auction must conclude with a Dutch auction when all bidders’ valuations are known to be below said price. A format such as beginning with an English auction at the monopoly price and then, should there be no takers, concluding with a Dutch auction is an example of an auction that is weakly shill-proof and efficient given the monopoly price. [^12] In fact, this auction format closely resembles the Honolulu–Sydney fish auction documented by Hafalir et al. (2023).
Weakly Shill-Proof and Ex-Post Incentive Compatible Optimal Auctions. The previous two results imply that ex-post incentive compatible auctions cannot be both strongly shill-proof and optimal, nor can they be (robustly) weakly shill-proof and efficient. The English auction (with an optimal reserve price) is, as we’ve noted, weakly shill-proof, optimal, and ex-post incentive compatible in the regular, IPV case. Is it the unique such auction? Does this combination of properties require a potentially large number of rounds? Our next result (Proposition 4.12) shows that, in general, the answer is no: In fact, although we can never find a weakly shill-proof, optimal, and ex-post incentive compatible auction that finishes in one round, we can always find a weakly shill-proof auction and a valuation distribution such that the worst-case number of auction rounds required is an arbitrarily small fraction of the number of rounds needed in the English auction.
Affiliated and Interdependent Values. Our last theorem, Theorem 5.4, shows that as the type distribution becomes less affiliated (Definition 5.2) and the values become less commonly valued (Definition 5.3), then both the set of strongly shill-proof auctions and the set of weakly shill-proof auctions expand. We then apply Theorem 5.4 to show most of our characterization results (Theorem 3.6, Theorem 4.3, and Corollary 4.8) hold in a more general, affiliated environment. Theorem 5.4 means that shill-proofness admits a partial order with respect to both the value function and the affiliation structure. Our theorem holds for any affiliation structure, most commonly used interdependent value functions (the function must satisfy Assumption 5.1), and any extensive-form auction where an optimal transfer rule is used. The last restriction is necessary because we consider discrete types and so there are a multitude of incentive compatible transfers that could yield different revenue. We generalize Lemma 3.1 in Proposition 5.7 to show that even with affiliation, a strongly shill-proof and optimal auction must be pay-as-bid with a reserve structure on the allocation.
@@
Proof (Theorem 4.3) (Proposition 4.12)
@@
Proof (Theorem 4.3) (Proposition 4.6)
@@
Proof (Proposition 4.6) (Robustly Unique, Proposition 4.6)
@@
Proof (Proposition 4.6) (Proposition 4.6)
@@
Proof (Proposition 4.6) (Unique, Proposition 4.6)
@@ 1.3 Related Work
While the idea and practice of shill bidding by a seller have long been well known, the auction theory literature on the topic is surprisingly thin. Chakraborty and Kosmopoulou (2004) consider common value auctions and focus on technological barriers (as opposed to auction formats) that can mitigate shill bidding. Lamy (2009) studies shill bidding specifically in English auctions in which bidders’ valuations are affiliated in the sense of Milgrom and Weber (1982), and proves that shill bidding effectively cancels out the effects of affiliation in equilibrium due to real bidders conditioning on bids being fake (see also Izmalkov (2004)). We also consider shill bidding in affiliated environments, but consider a much larger class of possible extensive-form games.
@@
Porter and Shoham (2005) consider a model similar to a second-price auction, motivated by “cheating” by online platforms that can announce a manipulated auction outcome subsequent to collecting all of the bidders’ bids. More recently, a number of works (e.g., Roughgarden (2021); Lavi et al. (2022); Basu et al. (2023); Chung and Shi (2023)) have considered shill bidding in the context of blockchain transaction fee mechanism design, with an emphasis on knapsack auctions that are ex-post incentive compatible, shill-proof, and robust to various forms of collusion. Ausubel and Milgrom (2006) and Day and Milgrom (2008) consider shill bids by bidders in a multi-item auction, who are looking to exploit complementarities to lower their payments in VCG-type mechanisms—as opposed to shill bids by a seller looking to increase revenue, as is the case of this paper. [^13] More broadly there is a literature on sybil-resistance referred to as “false-name proofness” (see, e.g., Conitzer et al. (2010) for a reference). Contemporaneous work by Shinozaki (2024) and Zeng (2024) also study shill bidding, but take a different approach to the problem. While we study ex-interim deterrence against shill bidding in extensive form games, they group auctions into equivalence classes based on the outcome and primarily focus on ex-ante deterrence against the seller inserting additional bidders into the auction to increase perceived competition. The results in those papers are complementary to ours: While we show that the dynamics of the auction are important for preventing shill bidding, they show that when the shill bidders can increase perceived competition outside of taking actions in the auction, only the posted-price mechanism is non-manipulable (see also Footnote 16).
@@
Our theory of shill-proof auctions is similar in spirit to the theory of credible mechanisms developed by Akbarpour and Li (2020), and leverages their framework for extensive-form auction analysis. That said, shill-proofness differs conceptually from credibility as shill-proofness focuses on the auctioneer’s incentives to insert fake bids, whereas credibility focuses on the auctioneer’s incentive to truthfully report the actions of a bidder to other bidders. Furthermore, the results in this paper are also qualitatively different. For example, there are a multitude of credible auctions, but only one strongly shill-proof auction and there are a multitude of strategy-proof and weakly shill-proof auctions, but only one strategy-proof and credible auction (see Section 6.1 for more discussion). In Proposition 6.1, we prove that strong shill-proofness is a stronger condition than credibility and weak shill-proofness is a weaker condition in the sense that if an auction is strongly shill-proof, then it is credible, which in turn implies that it is weakly shill-proof.